OpenStudy (anonymous):

Please Help! Use the first three non-zero terms of the maclaurin series to approximate the value of the definite interval from 0 to 0.8 (1-cost)/t^2 dt

OpenStudy (anonymous):

\[\int\limits_{0}^{0.8} (1-cost)/t^2 \]

OpenStudy (anonymous):

sint = x - x^3/3! + x^5/5! cost = 1 - x^2/2! + x^4/4! 1 - cost = t^2/2! - t^4/4! substitute and find..

OpenStudy (anonymous):

I got to that part, then I divided by 2 to get 1/2 - t^2/24+x^4/720 ... then I don't know what to do